# Tree balance in phylogenetic models

### v2

@@ 4 Results
This section contains the most interesting results, i.e., only a selection of alternative tree models, of the simulation studies described in Section 3. The complete collection of result figures can be found in Section 4 of the supplementary material [27].
@@
In brief, all figures show the power (with 95%-confidence bands) of the various TSS to correctly identify trees generated under models distinct from the Yule model as not having been generated under the Yule model. The power is plotted either against the leaf number $n$ (Figure 3) or against the model’s parameters (Figures 49).
Comparing the different figures, several observations spring to mind. First, and maybe unsurprisingly, the curves look very different for different models. In some cases, there are few to no crossings between the power curves for different TSS (e.g., Figures 4 and 7), implying that the ranking of the TSS by their power does not change across the different settings tested, whereas for others, there are significant changes (e.g., Figures 3, 6, or 8). We also note that symmetric (parent species splits into two child species) versus asymmetric speciation (one child species splits off but parent species persists) significantly impact the results (see, e.g., Figure 8).
@@
Second, there are some TSS that (almost) consistently outperform the others, such as the $\widehat{s}$-shape statistic, whereas others, such as the modified maximum difference in width, tend to perform rather poorly. In both cases, there are notable exceptions, though. In case of the $\widehat{s}$-shape statistic it is interesting to note that it is outperformed by other TSS (such as the $B_{2}$ index, the Colless index, the Sackin index, and the Mean $I^{\prime}$ index) for parameter-model combinations that are close to the Yule model (see Figures 6, 8). In case of the modified maximum difference in widths, exceptions to its overall relatively poor performance are for instance Aldous’ $\beta$ splitting model (Figure 4) and the biased speciation model (Figure 6). Moreover, while the power of TSS tends to increase with increasing $n$ (see Figure  3 and the differences between $n=30$ and $n=100$ in Figure 8), the power of the modified maximum difference in widths does not seem to follow this pattern.
@@
Third, sometimes several TSS perform almost equally well or equally poorly (indicated by their curves being clumped together as it is for instance the case for the top three curves in Figure 7), whereas in other cases the curves are clearly separated (e.g., Figure 6). Interestingly, the total cophenetic index as well as the quadratic Colless index show extremely similar behavior (their lines are nearly always right on top of each other), and it would be interesting to investigate this further.
While there are many more observations to be made by carefully studying the different plots, these observations already show that there are significant performance differences between the various TSS, which need to be considered when TSS are used in the context of macroevolutionary models.
@@
Figure 3: This figure displays the power of all TSS to correctly identify trees generated under the PDA model as not having been generated under the Yule model for several choices of $n$.
@@
Figure 4: This figure displays the power of all TSS to correctly identify trees generated under the Aldous’ $\beta$ splitting model as not having been generated under the Yule model. Since for $\beta=0$, Aldous’ splitting model corresponds to the Yule model, all TSS rejected $\approx 0.05$ of the trees in that case (as specified with the level of significance).
Figure 5: This figure displays the power of all TSS to correctly identify trees generated under Ford’s $\alpha$ model as not having been generated under the Yule model. Since for $\alpha=0$, Ford’s $\alpha$ model corresponds to the Yule model, all TSS rejected $\approx 0.05$ of the trees in that case (as specified with the level of significance).
Figure 6: This figure displays the power of all TSS to correctly identify trees generated under the biased speciation model as not having been generated under the Yule model.
@@
Figure 7: This figure displays the power of all TSS to correctly identify trees generated under the asymmetric simple Brownian model as not having been generated under the Yule model.
@@
Figure 8: This figure displays the power of all TSS to correctly identify trees generated under the (a)symmetric punctuated(-intermittent) log-Brownian model as not having been generated under the Yule model. The starting trait value is $x_{0}=10$ with constant $\sigma_{x}=1$.
@@
Figure 9: This figure displays the power of all TSS to correctly identify trees generated under the alternative birth-death model as not having been generated under the Yule model.
5 Software
We now present our software package poweRbal, implemented in the free programming language $\mathsf{R}$, which provides the readers with the means to do similar analyses as shown in the preceding section and adapt them to their needs. The package will shortly be published on CRAN. Using this package, the user can freely specify the set of tree shape statistics (all tree balance indices implemented in the package treebalance [13] are already included and can be used, but additional TSS can be added by the user as well), the null model, the set of alternative models (all tree models analyzed in this manuscript have already been included in the package and can be used, but additional models can be added by the user as well), the number of leaves $n$, and the method parameters like the sample size $N_{d}$ or the level of significance.
Below, we give a short introduction on how to use the package using illustrative examples. First, the package is loaded with library("poweRbal").
poweRbal provides the means to generate trees under various tree models. For example, if we want to generate a tree specifically under the PDA model, we can use genPDATree(n) (information on the function names of the individual tree models is given in the tables in Section 3 of the supplementary material [27]). Another way to access any tree model and produce any number of trees is to use the wrapper function genTrees(n, Ntrees, tm). With n we specify the number of leaves, with Ntrees the number of trees, and with tm the tree model either as a string/name if the tree model does not have or need a parameter or as a list of a string/name and parameter (see ?genTrees for more information on how to access the individual models). For example, we can use the following commands to generate 5 trees with 10 leaves each, once under the Yule model (default: rate 1), once under the PDA model, and once under Aldous’ $\beta$ splitting model with $\beta=-1$:
genTrees(n = 10, Ntrees = 5, tm = "yule")
genTrees(n = 10, Ntrees = 5, tm = "pda")
genTrees(n = 10, Ntrees = 5, tm = list("aldous", -1))
To compare the power of pre-implemented TSS under pre-implemented tree models, we can use the function powerCompare (see ?powerComp for detailed information). For this, we only have to specify the TSS, the alternative models, as well as the number of leaves. Optionally, the user can also specify the null model (by default the Yule model) and several other options that define the method (by default: sample size = 1,000, level of significance = 0.05, two-tailed test). For example, to use the Sackin and $B_{1}$ index (see tssInfo for information on how to access all available TSS) for the PDA and ETM models, we specify:
pc1 <- powerComp(tss = c("Sackin", "B1I"), alt_models = list("pda", "etm"), n = 10)
# Which yields the following matrix:
pda etm
Sackin 0.255 0.129
B1I 0.213 0.106
As another example, to use the Sackin and $B_{1}$ index for Aldous’ $\beta$ splitting model with $\beta=-1.5$, $\beta=-1$, and $\beta=-0.5$, we have:
pc2 <- powerComp(tss = c("Sackin", "B1I"), n = 10,
alt_models = list(list("aldous", -1.5), list("aldous", -1),
list("aldous", -0.5)))
# Which yields the following matrix:
Sackin 0.272 0.087 0.043
B1I 0.242 0.071 0.044
The first results can be plotted [^2] The figures in Section 4 have been created similarly, but with the help of the ggplot2 package for aesthetic reasons. with showPower(power_data = pc1) and the second with
showPower(power_data = pc2) or – better for representatives of parametrized tree model families as in this case – with showPower_param(power_data = pc2, model_param = c(-1.5, -1, -0.5), ylim = c(0,0.4)) (see the help files accessible with ?showPower and ?showPower_param for more information on customizing these plots).
To compare new TSS (the example below combines the Sackin index and the total $I^{\prime}$ index from [25, p.53] into a new TSS) with the existing ones, we can create a list object containing the function and a short name (and optionally more information as in tssInfo).
library("treebalance")
new_tss_func <- function(tree){
n <- length(tree\$tip.label)
return(sackinI(tree)+ 0.14*n*IbasedI(tree, method = "total",
correction = "prime", logs = F))
new_tss <- list(func = new_tss_func, short = "SackTotI")
Similarly, we can add new tree models by creating a list object containing the function which produces Ntrees-many trees with $n$ leaves. With that, it is also possible to use lists of trees, e.g., from a tree model that was implemented in a different programming language, as a basis for a null or alternative model. The ape package [35] provides means for importing common file formats like the Newick format. As an example for this, let treeList be a list of type multiphylo containing trees with the same number of leaves, say $n=30$. Then we could proceed as follows.
new_tm_func <- function(n , Ntrees = 1){
if (n == 30) {
selection <- sample.int(length(treeList), Ntrees)
return(treeList[selection])
} else {
stop("No such trees available.")
new_tm <- list(func = new_tm_func)
Then, we add the names of the list objects as follows. Here, the new tree model is used as an alternative model but it can be also used as the null model by setting null_model = "new_tm".
pc3 <- powerComp(tss = c("Sackin", "new_tss", "B1I"),
alt_models = list("pda", "new_tm"), n = 10)
showPower(power_data = pc3)
@@ 6 Conclusion and outlook
In most of our analyses, the $\widehat{s}$-shape statistic was the most powerful statistic discriminating alternative models from the Yule model. While this nicely complements a study by Blum at al. [8], who found that the $\widehat{s}$-shape statistic “warrants maximal power for rejecting the [Yule] model against the PDA [model]”, this is hardly surprising. It can easily be explained by the fact that the $\widehat{s}$-shape statistic was even introduced with that very purpose in mind. In fact, except for the normalizing constant of $P_{Y,n}$, the $\widehat{s}$-shape statistic corresponds to the logarithm of $P_{Y,n}$ (c.f. [46]), i.e., of the probability of generating a tree under the Yule model. However, while Blum et al. in their study explicitly did not claim that the $\widehat{s}$-shape statistic is “generally superior”, in our study we could see that it is indeed superior concerning many models, not only the PDA model. Some exceptions include the Brownian models, in which the $\widehat{s}$-shape statistic is outperformed by the $B_{2}$, Sackin and Colless indices, c.f. Figures 7 and 8. Khurana et al.  [28] also found that for their data, the $\widehat{s}$-shape statistic’s performance was mediocre, while the $B_{2}$ index, the variance of leaf depths, $I$-based indices, stairs2, and the normalized Colless index were most powerful in discriminating empirical trees from trees simulated under a constant-rate-birth-death model (which induces the same distribution on $\mathcal{BT}^{\ast}_{n}$ as the Yule model). Their findings concerning $B_{2}$ nicely complement a study by Bienvenu et al. [6], which used five different null models (one of which was the Yule model) and tested six different balance indices for their power (note that they did not include the $\widehat{s}$-shape statistic in their study, though). In this study, the authors found that the $B_{2}$-index is on average better than the other investigated indices. Note that in most of our studies, too, the $B_{2}$ index was amongst the most powerful indices.
As mentioned above, the choice of different null models can make a huge difference. However, even with the Yule model as a null model, there are huge performance differences between the different TSS. These differences might imply that there are different patterns of imbalance produced by the different models, some of which can only be recognized by a subset of tree shape statistics. Therefore, the choice of the most powerful TSS for each individual case is crucial.
However, while we have shown performance differences between various balance indices, most of them (interestingly, with the exception of maximum width over maximum depth) have in common that generally their power increases with increasing $n$, i.e., with an increasing number of species under investigation. This information might be helpful for biologists in order to plan their studies.
Due to limited space in this manuscript, we have only listed a few example comparisons here. However, a comprehensive list of a wider collection of TSS / model combinations induced by our manuscript can be found in the supplementary material [27].
@@
We are confident that our manuscript will inspire future research. Concerning applications, our software package poweRbal is specifically designed in such a way that it also caters to researchers unfamiliar with the programming language R. Trees generated by other software tools can easily be loaded into our package (e.g., in the well-known Newick format) and used as a basis for an alternative model. Newly developed indices and models can easily be added to the package, so that the user can test several variations to see which one is most powerful in the given setting. poweRbal also enables researchers to perform a simulation study beforehand to decide on which TSS to use and to subsequently analyze their data with the respective optimal tools.
@@
Finally, our manuscript has focused on basic models, i.e., on models which contain clear instructions on how to generate (phylogenetic) trees. However, in the literature also more elaborate and complex models can be found (e.g., [32, pp. 9–13] and [30]). For instance, evolution can be modeled at the genome level, so that subsequently a phylogenetic tree can be reconstructed (e.g., [30]). For the latter, different reconstruction methods can be used, and it has been shown that these methods also have an impact on the resulting tree’s balance (see, e.g., [40, 23], [9, p. 1141])). Moreover, we have primarily focused on models from phylogenetics, but tree models also frequently occur in other areas such as population genetics, probability theory, or computer science. While some of the models used in these areas are closely related to the Yule model (see, e.g., [14] and Remark 3.1 in the supplement of the present manuscript), it would be an interesting direction of future research to include models such as Aldous’ continuum random tree model [2, 3, 4] in the analyses.
Acknowledgement.
MF and SK were supported by the project ArtIGROW, which is a part of the WIR!-Alliance “ArtIFARM – Artificial Intelligence in Farming”, and gratefully acknowledge the Bundesministerium für Bildung und Forschung (Federal Ministry of Education and Research, FKZ: 03WIR4805) for financial support. Moreover, we wish to thank Volkmar Liebscher for helpful advice on the development of the R-package and Annalena Franke for implementing first versions of several tree models.

### v3

@@ 4 Results
In this section, we summarize the most important results obtained in our simulation study. Figures 37 show the power (with 95%-confidence bands) of the various TSS to correctly identify trees generated under models distinct from the Yule model as not having been generated under the Yule model. The power is plotted either against the leaf number $n$ (Figure 3) or against a tree model parameter (Figures 47). Note that all figures follow the legend of Figure 3. Further result figures can be found in Section 5 of the supplementary material [40]. As evident from all our figures, the curves look quite different for different alternative models. In the following, we study some aspects of these figures and conclusions to be drawn more in-depth.
“Good” and “bad” TSS
A central goal of our study was to identify TSS that exhibit a high power for discriminating between different (macro)evolutionary models. Our results show that that the performance of different TSS clearly depends on the alternative models and also on the null model used (recall that we use the Yule model as the null model throughout the present manuscript, but in Section 6 of the supplementary material [40] we also briefly discuss other null models). Even for a fixed alternative model, there are sometimes striking differences for different parameter choices. For example, considering the log-Brownian models ins Figure 6, we can observe that the $B_{2}$ index and the equal weights Colless index are among the worst TSS for small values of $\sigma_{\lambda}$, whereas they are among the most powerful TSS for larger values of $\sigma_{\lambda}$. A related observation is that in some cases, there are few to no crossings between the power curves for different TSS (e.g., Figures 4 and 5), implying that the ranking of the TSS by their power does not change across the different settings tested, whereas for others, there are significant changes (e.g., Figures 5 and 6).
@@
Nevertheless, we can observe that there are some TSS that (almost) consistently outperform the others, such as the $\widehat{s}$-shape statistic, whereas others, such as the modified maximum difference in width or the rooted quartet index, tend to perform rather poorly. In all cases, there are notable exceptions, though. In the case of the $\widehat{s}$-shape statistic it is interesting to note that it is outperformed by other TSS (such as the $B_{2}$ index or the Sackin index) in the symmetric simple Brownian model (Figure 5) as well as for parameter-model combinations close to the Yule model in the asymmetric log-Brownian model (Figure 6)). In case of the modified maximum difference in widths, exceptions to its overall relatively poor performance are for instance Aldous’ $\beta$ splitting model (Figure 4) and the biased speciation model (Figure 5). Interestingly, the biased speciation model is also an exception to its overall relatively poor performance. Another exception for the rooted quartet index is the simple Brownian model (Figure 5). It is also noticeable that while the power of TSS should increase with increasing $n$, and this is indeed the case for almost all TSS (see Figure  3 and the differences between $n=30$ and $n=100$ in Figure 6), the power of the modified maximum difference in widths does not seem to follow this pattern for the PDA model (Figure 3).
Additionally, it is interesting to note that Fischer et al. [78, Chapter 26] recently observed that the $\widehat{s}$-shape statistic had a high resolution for two empirical data sets (meaning that it showed few ties), whereas the modified maximum difference in widths showed a low resolution (meaning that it showed many ties). This could indicate a connection between resolution and power for some but not all TSS. For the rooted quartet index, for instance, Fischer et al. [78, Chapter 26] observed a high resolution while we observe a relatively low power of the rooted quartet index in the present analysis. Nevertheless, exploring why certain TSS behave in distinct ways and whether there is a connection between power and resolution for some TSS is an interesting direction for future research.
@@
Finally, we remark that it is often the case that several TSS perform almost equally well or equally poorly, indicated by their curves being clumped together, as is for instance the case for the top curves in the simple Brownian model (Figure 5). In other cases, the curves are clearly separated, as for instance in the biased speciation model (also Figure 5). Interestingly, the total cophenetic index and the quadratic Colless index show extremely similar behavior (their lines are nearly always right on top of each other), and it would be interesting to investigate this further. In summary, we thus reiterate that what is the best and worst TSS depends on the model setting, highlighting the importance of performing a power analysis such as the one presented here prior to downstream analyses.
Fitting tree models
For parameter settings that correspond to the Yule model, the power of all TSS should be close to 0.05 and this is clearly visible in the corresponding figures (e.g., for $\beta=0$ or $\alpha=0$ in Figure 4, or $\mu=0$ in Figure 7). Interestingly, we also observe a power close to 0.05 for almost all TSS for the biased speciation model for $\zeta$ between 0.20 and 0.25 (Figure 5). This could indicate that the biased speciation model is very close to or coincides with the Yule model for $\zeta$ in this range. However, it could also mean that none of the TSS employed in this study can detect deviation from the Yule model in this setting, even though the models might theoretically not be identical. This leads us to the following observation for fitting tree models to empirical data: When a collection of empirical trees is tested against a specific evolutionary model and a low power is observed for all TSS, this can be an indication of a good fit of the model to the data. However, low power is only a sufficient and not a necessary condition for a good model fit, as it could be the case that other TSS are needed to detect differences between the data and the model. Thus, caution is warranted. A good example illustrating this can be found in Figure 20 in the supplementary material [40], where we use Aldous’ $\beta$ splitting model with $\beta=-1$ as the null model and the alternative birth-death model as the alternative model. In this case, almost none of the TSS can tell the alternative birth-death trees with parameter $\mu=0.4$ apart from the Aldous’ $\beta=-1$ trees, except for the $B_{2}$ index and the Furnas rank (where the latter is even more surprising as it is primarily used as a tree enumeration method and not often employed in applications). Without including these two indices, one might have concluded that both models are equivalent.
Influence of the speciation mode
[…]
@@
Figure 3: The power of all TSS to correctly identify trees generated under the PDA model as not having been generated under the Yule model as a function of $n$. Note that for $n=10$ and $n=15$ the power values are based on the exact TSS distribution under Yule.
@@
Figure 4: The power of all TSS to correctly identify trees generated under the Aldous’ $\beta$ splitting model (left) and Ford’s $\alpha$ model (right) as not having been generated under the Yule model as a function of $\beta$ (resp. $\alpha$). Since for $\beta=0$, Aldous’ splitting model corresponds to the Yule model, all TSS rejected $\approx 0.05$ of the trees in that case (as specified with the level of significance). Similarly, since for $\alpha=0$, Ford’s $\alpha$ model corresponds to the Yule model, all TSS rejected $\approx 0.05$ of the trees in that case.
@@
Figure 5: The power of all TSS to correctly identify trees generated under the biased speciation model (left) and the asymmetric simple Brownian model (right) as not having been generated under the Yule model as a function of $\zeta$ (resp. $\sigma$).
@@
Figure 6: The power of all TSS to correctly identify trees generated under the symmetric (left) and asymmetric (right) punctuated(-intermittent) log-Brownian model as not having been generated under the Yule model as a function of $\sigma_{\lambda}$. The starting trait value is $x_{0}=10$ with constant $\sigma_{x}=1$.
@@
Figure 7: The power of all TSS to correctly identify trees generated under the alternative birth-death model as not having been generated under the Yule model as a function of $\mu$ for two different different tree sizes ($n=30$ and $n=100$).
@@ 5 Conclusion and outlook
In most of our analyses, the $\widehat{s}$-shape statistic was the most powerful statistic discriminating alternative models from the Yule model. While this nicely complements a study by Blum at al. [27], who found that the $\widehat{s}$-shape statistic “warrants maximal power for rejecting the [Yule] model against the PDA [model]”, this is hardly surprising. It can easily be explained by the fact that the $\widehat{s}$-shape statistic was even introduced with that very purpose in mind. In fact, except for the normalizing constant of $P_{Y,n}$, the $\widehat{s}$-shape statistic corresponds to the logarithm of $P_{Y,n}$ (c.f. [79]), i.e., of the probability of generating a tree under the Yule model. However, while Blum et al. in their study explicitly did not claim that the $\widehat{s}$-shape statistic is “generally superior”, in our study we could see that it is indeed superior concerning many models, not only the PDA model. Some exceptions include the Brownian models, in which the $\widehat{s}$-shape statistic is outperformed by other TSS, c.f. Figures 5 and 6. Khurana et al.  [42] also found that for their data, the $\widehat{s}$-shape statistic’s performance was mediocre, while the $B_{2}$ index, the variance of leaf depths, $I$-based indices, stairs2, and the normalized Colless index were most powerful in discriminating empirical trees from trees simulated under a constant-rate-birth-death model (which induces the same distribution on $\mathcal{BT}^{\ast}_{n}$ as the Yule model). In Section 6 of the supplementary material we discuss their findings further and show how to approach their research questions with our methods. We also remark that the findings by Khurana et al.  [42] concerning $B_{2}$ nicely complement a study by Bienvenu et al. [80], which used five different null models (one of which was the Yule model) and tested six different balance indices for their power (note that they did not include the $\widehat{s}$-shape statistic in their study, though). In this study, the authors found that the $B_{2}$-index is on average better than the other investigated indices. Note that in most of our studies, too, the $B_{2}$ index was amongst the most powerful indices.
@@
As shown by Bienvenu et al. [80], the choice of different null models can make a huge difference. However, as our results show, even with the Yule model as a null model, there are huge performance differences between the different TSS. These differences might imply that there are different patterns of imbalance produced by the different models, some of which can only be recognized by a subset of tree shape statistics. Therefore, the choice of the most powerful TSS for each individual case is important. We thus highly recommend to perform a power analysis such as the one presented here prior to downstream analyses. For this purpose, our software package poweRbal is specifically designed in such a way that it also caters to researchers unfamiliar with the programming language R. Trees generated by other software tools can easily be loaded into our package (e.g., in the well-known Newick format) and used as a basis for an alternative model. Newly developed indices and models can easily be added to the package, so that the user can test several variations to see which one is most powerful in the given setting. poweRbal also enables researchers to perform a simulation study beforehand to decide on which TSS to use and to subsequently analyze their data with the respective optimal tools.
@@
Finally, our manuscript has focused on basic models, i.e., on models which contain clear instructions on how to generate (phylogenetic) trees. However, in the literature also more elaborate and complex models can be found (e.g., [41, pp. 9–13] and [81]). For instance, evolution can be modeled at the genome level, so that subsequently a phylogenetic tree can be reconstructed (e.g., [81]). For the latter, different reconstruction methods can be used, and it has been shown that these methods also have an impact on the resulting tree’s balance (see, e.g., [82, 83], [84, p. 1141])). Moreover, we have primarily focused on models from phylogenetics, but tree models also frequently occur in other areas such as population genetics, probability theory, or computer science. While some of the models used in these areas are closely related to the Yule model (see, e.g., [64] and Remark 3.1 in the supplement of the present manuscript), it would be an interesting direction of future research to include models such as Aldous’ continuum random trees  [85, 86, 87] in the analyses.
Acknowledgement
MF and SK were supported by the project ArtIGROW, which is a part of the WIR!-Alliance “ArtIFARM – Artificial Intelligence in Farming”, and gratefully acknowledge the Bundesministerium für Bildung und Forschung (Federal Ministry of Education and Research, FKZ: 03WIR4805) for financial support. Moreover, we wish to thank Volkmar Liebscher for helpful advice on the development of the R-package and Annalena Franke for implementing first versions of several tree models. Additionally, we thank three reviewers and the editor Noah Rosenberg for detailed comments on an earlier version of this paper.